A cracker rocket is ejecting 80gm gases per minute at the speed of 420m/s.Find the force on the rocket?

Grade:11

1 Answers

Samyak
6 years ago
Force=maM=80gm/min=80/(1000*60)kg/sNow force=vdm/dt=420*80/(60000)=42/75N/sThat is it is acted by this force every second.